If log 2 » 0.301 and log 3 » 0.477, then log 6 »
Found 2 solutions by Earlsdon, stanbon:
Answer by Earlsdon(6294)

Try this:
Log(6) = Log(2*3) = Log(2)+log(3) = 0.301+-.477 = 0.778

Answer by stanbon(75887)

If log 2 » 0.301 and log 3 » 0.477, then log 6
--------------------
log6 = log(2*3) = log2 + log3 = 0.301+0.477 = 0.778151
==================
Cheers,
Stan H.